OpenStudy (anonymous):

Solve √x+6+√x=3 for x. A)1/4 B) 1/2 C)2 D)2 --thanks :D

OpenStudy (anonymous):

hey is 6 also in square root or just outside it?

OpenStudy (anonymous):

@heybhai its also in the square root :)

OpenStudy (anonymous):

think the best thing to do is to substitute the values, since it's a MCQ

OpenStudy (anonymous):

@kelvinltr Lol whut. ._. I'm slow when it comes to math lol sorry

OpenStudy (anonymous):

ya @kelvinltr is right just put 1/4 is eqn1

OpenStudy (anonymous):

i mean, try plugging in x=1/4, x=1/2 etc into the equation and see whether you get the left hand side and right hand side equal, rather than trying to solve it

OpenStudy (anonymous):

as u put sqrt(1/4+6)+sqrt(1/4) sqrt(25/4)+sqrt(1/4) 5/2+1/2(taking LCM) 6/2 =3

OpenStudy (anonymous):

Ohhh okay! :D

OpenStudy (anonymous):

option 1 is answer

OpenStudy (anonymous):

@Heybhai yeah I just got that answer :3 Thanks you guys :D